Ripley Corporation has supplied the following information obtained from its standard cost system in June.

Standard price of direct materials $ 6 per pound
Actual price of direct materials $ 5 per pound
Standard direct labor rate $ 9 per hour
Actual direct labor hours in June 9,500 hours

The following journal entries were made during June with respect to Ripley’s standard cost system.

General Journal Debit Credit
Work in Process Inventory (at standard cost) $ 48,000
Materials Quantity Variance 1,200
Direct Materials Inventory (at actual cost) $ 41,000
Materials Price Variance 8,200
To record the cost of direct materials used in June.
Work in Process Inventory (at standard cost) $ 81,000
Labor Rate Variance 950
Labor Efficiency Variance 4,500
Direct Labor (at actual cost) $ 86,450
To record the cost of direct labor charged to production in June.
Work in Process Inventory (at standard cost) $ 25,000
Overhead Spending Variance 2,000
Overhead Volume Variance $ 5,000
Manufacturing Overhead (at actual cost) 22,000
To apply overhead to production in June.

Required:

a. Determine the actual quantity of materials purchased and used in production during June.

b. Determine the standard quantity of materials allowed for the productive output achieved during June.

c. Determine the actual average direct labor rate in June.

d. Determine the standard direct labor hours allowed for the production output achieved during June.

e. Determine the total overhead costs allowed for the production output achieved during June.

f. Prepare a journal entry to record the transfer of all work in process to finished goods at the end of June.

g. Close all cost variances directly to the Cost of Goods Sold account at the end of June.

h. Was Ripley's production output in June more or less than its normal level of output?
Answer #1

(a) Material price variance = Actual quantity used * (Standard price - Actual price)

$8,200 = Actual quantity used * ($6 - $5)

Actual quantity used = $8,200 * $1 = $8,200 pounds

(b) Material quantity variance = Standard price * (Standard quantity - Actual quantity)

($1,200) = $6 (Standard quantity - 8,200)

($1,200) = $6 (Standard quantity) - 49,200

Standard quantity = $48,000 / $6 = $8,000 pounds

(c) Labor rate variance = Actual labor hours * (Standard rate - Actual rate)

(950) = 9,500 * ($9 - Actual rate)

(950) = 85,500 * 9,500(Actual rate)

Actual rate = $86,450 / 9,500 = $9.10 per hour

(d) Labor efficiency variance = Standard rate * (standard hours - actual hours)

(4,500) = $9 * ( Standard hours - 9500)

(4,500) =$9SH - 85500

$9 Standard hours = $81,000 / 9= 9000 hours

e)

Total overhead costs = 22000 - 2000 = $20000

f)

Debit Credit
Finished goods inventory 154000
work in process inventory 154000

Please post other sub parts separately. As per HomeworkLib policy, more than 4 sub parts are solved